OpenStudy (anonymous):

Use the perfect square trinomial formulas to factor, factor out any common factors. 16y^2 - 8y + 1

OpenStudy (anonymous):

Split (16y^2 -8y) from +1, and you'll get (8y)(2y-1)+1. Or by the quadratic equation: \[[8\pm \sqrt{8^2 -4(16)(1)} ] /2(16) = (8\pm0)/16 = 1/2, then, (x+1/2) is the only \root\]

OpenStudy (anonymous):

By the quadratic (formula above is messed up), the only root is (x+1/2)